Lourdes is decorating a toy box for her sister. She will use self-adhesive paper to cover all of the exterior sides except for the bottom of the box. The toy box is 4 feet long, 3 feet wide, and 2 feet high. How many square feet of adhesive paper will Lourdes use to cover the box?
Answer by mananth(16946)

Lourdes is decorating a toy box for her sister. She will use self-adhesive paper to cover all of the exterior sides except for the bottom of the box.
The toy box is 4 feet long, 3 feet wide, and 2 feet high. How many square feet of adhesive paper will Lourdes use to cover the box?
4 feet long,L
3 feet wide,W
and 2 feet high.H
The surface area of the box
2LW + 2WH + 2LH
base not covered
so tape required = LW +2WH + 2LH

= 4*3 +2*3*2+2*4*2
=12+12+16
=40 ft^2
=